### v1

# FSDAM: Few-Shot Driver Attention Modeling via Vision-Language Coupling

Abstract
Understanding where drivers look and why they shift their attention is essential for autonomous systems that read human intent and justify their actions. Most existing models rely on large-scale gaze datasets to learn these patterns; however, such datasets are labor-intensive to collect and time-consuming to curate. We present FSDAM (Few-Shot Driver Attention Modeling), a framework that achieves joint attention prediction and caption generation with approximately 100 annotated examples, two orders of magnitude fewer than existing approaches. Our approach introduces a dual-pathway architecture where separate modules handle spatial prediction and caption generation while maintaining semantic consistency through cross-modal alignment. Despite minimal supervision, FSDAM achieves competitive performance on attention prediction, generates coherent, and context-aware explanations. The model demonstrates robust zero-shot generalization across multiple driving benchmarks. This work shows that effective attention-conditioned generation is achievable with limited supervision, opening new possibilities for practical deployment of explainable driver attention systems in data-constrained scenarios.
1 Introduction
Driving is primarily a visual task. Drivers constantly move their gaze across different environmental elements, such as traffic signals, crossing pedestrians, and road hazards. This selective attention directly influences their understanding of the situation and decision-making process. Suboptimal attention allocation due to distraction or drowsiness remains a leading cause of accidents, with disproportionate crash rates at intersections where drivers fail to notice crossing traffic, or during lane changes when blind spots go unchecked [9312486, 10588743]. Understanding where and what drivers visually attend to can also be beneficial to Advanced Driver Assistance Systems (ADAS) and autonomous vehicles (AVs) operating in mixed-traffic environments, where human-AI collaboration depends on a shared perception of the surroundings. Modeling driver attention allows for cross-verification of the human-machine goal alignment by comparing human intention with decisions made by autonomous systems. This process validates whether the planned maneuvers are consistent with the perceived context of the scene, and it also supports interpretable decision-making chains that enhance user trust in human-AI collaboration.
Figure 1: Few-shot learning performance on BDD-A dataset. Gaze prediction metrics: $C$, SIM, NSS, AUC-J, AUC-B, and KL (inverted). All metrics are linearly min-max normalized to [0,1] where larger area indicates better performance.
Despite its importance, effectively modeling driver attention faces fundamental challenges. Existing approaches rely heavily on supervised learning that requires large, labeled eye-tracking data collected in well-represented conditions to provide stable inductive biases for low-level perception. However, the data collection and annotation can be costly and often lack diversity in safety-critical scenarios [li2025recogdrive]. Additionally, an image dataset containing driver gaze information can expose sensitive biometric and behavioral data, potentially revealing individuals’ identity and attention patterns [Kröger2020]. This data scarcity, coupled with privacy and annotation costs, can lead to long-tail errors during domain-specific corner cases and rare events [ghosh2025roadwork]. Moreover, the space of possible driving scenarios grows combinatorially with environmental factors, traffic configurations, and agent behaviors, making comprehensive coverage through dataset expansion increasingly impractical [ghosh2025roadwork, li2025recogdrive]. These constraints necessitate a shift from data-hungry supervised learning to sample-efficient learning with semantic grounding. In addition to learning visual attention patterns through labeled datasets, semantic understanding of the surrounding environment can also benefit visual attention prediction. For example, knowing that “pedestrians near crosswalks" or “vehicles braking suddenly" are situations where driver attention is typically expected, and can be encoded as an auxiliary loss that encourages the model’s attention distribution to place more weight on these regions. Vision-language models (VLMs) can offer this capability: their pre-trained representations encode rich semantic knowledge about objects, spatial relationships, and causal reasoning. Rather than learning attention patterns from scratch, VLMs can exploit these semantic priors to achieve few-shot generalization with as few as 5-32 examples per class [radford2021learning]. In autonomous driving, works like DriveLM [sima2023drivelm] have demonstrated that grounding perception in natural language enables more interpretable reasoning that generalizes across scenarios. Building on this insight, Zhou et al. [zhou2025where] pioneered joint attention-language modeling with LLaDA, demonstrating that coupling spatial attention prediction with textual reasoning improves both quantitative performance and interpretability. Similarly, Chen et al. [chen2024gazexplainlearningpredictnatural] showed that generating natural language explanations enhances gaze prediction understanding. While these works demonstrate the potential of vision-language models for joint spatial and semantic reasoning, they still require extensive labeling. The objective of this work is therefore to leverage VLM priors to learn driver attention from fewer labeled examples.
To address these limitations, we develop FSDAM, a framework that learns joint gaze prediction and attention-grounded caption generation from approximately 100 annotated examples. To our knowledge, this is the first work to demonstrate that attention-conditioned language generation can be learned effectively in a few-shot regime. Beyond data efficiency, we introduce attention shift anticipation: unlike prior work that describes current attention, our model predicts where drivers will look next from single frames.c We create supervision by selecting frame pairs where attention demonstrably shifts and annotating observed transitions with structured captions describing scene context, current focus, anticipated shift, and causal reasoning. At inference, the model learns to anticipate these shifts from single frames alone, capturing the forward-looking nature of human attention without requiring video input. Our key insight is to structure the learning task so that spatial attention and language understanding provide complementary supervision signals. We introduce a structured attention reasoning representation where each annotation describes the driving scene, identifies where the driver is currently looking, predicts where attention will shift next, and explains why. This structured representation creates explicit connections between visual attention patterns and semantic reasoning.
Our architecture processes visual features through two pathways: one generates attention-grounded captions while the other aligns predicted attention with language representations. Both pathways share the same visual features but specialize for their respective tasks. Starting from approximately 100 training examples curated from the Berkeley DeepDrive Attention (BDD-A) dataset [10.1007/978-3-030-20873-8_42], our model achieves strong spatial attention prediction despite the severe data limitation. It reduces KL divergence on BDD-A by up to 89% compared to fully supervised baselines and delivers competitive CC and NSS values. Zero shot transfer experiments on DR(eye)VE [palazzi2018predicting] and DADA 2000 [fang2023dadadriverattentionprediction] further show that FSDAM maintains the lowest KL divergence and strong SIM scores across these unseen domains, indicating that the learned attention language alignment generalizes well beyond its training distribution. This efficient learning regime supports practical use in driving environments where gaze data is limited, such as night scenes, weather affected roads, and high speed highways. In summary, we make the following contributions:
  • First few-shot approach for attention-based generation. We are the first of its kind to achieve joint spatial attention prediction and natural language explanation in a few-shot learning regime, training from fewer than 100 examples, two orders of magnitude more data efficient than existing joint modeling approaches [zhou2025where, chen2024gazexplainlearningpredictnatural].
  • Attention shift anticipation. Instead of focusing on current attention, our model introduces a structured attention reasoning representation that captures scene context, current focus, anticipated shift and causal factors. This enables models to predict where attention will move next from single frames, capturing the forward-looking nature of human attention without requiring video input at inference.
  • Dual-pathway architecture for joint learning. We design an architecture with separate pathways for caption generation and attention-language alignment that share visual representations, enabling task-specific specialization while maintaining consistency.
  • Strong empirical validation. Despite training in few-shot regime, we achieve competitive gaze prediction and caption generation with strong zero-shot transfer across three standard benchmarks demonstrating generalization beyond the training data distribution.
2 Related Works
2.1 Driver Attention Modeling
Driver attention modeling has been studied as a visual saliency prediction task, estimating spatial gaze heatmaps from dashcam images or video sequences [zhou2025where]. Early works applied CNNs to learn correlations between scene features and eye fixations, while subsequent models incorporated temporal dynamics and multi-modal inputs using CNN-LSTMs and feature fusion from RGB, optical flow, and semantic segmentation [7789504, 10.1007/978-3-030-20873-8_42]. The SEEV model [steelman2017theory] outlines bottom-up (salience, effort) and top-down (expectancy, value) factors influencing attention allocation. Existing models typically emphasize either low-level visual features or task-level signals such as GPS and traffic semantics [8751968, 10588743]. These approaches provide limited insights into the underlying causes of attention allocation beyond saliency.
2.2 Vision-Language Models for Explainable Attention
Recent work has applied VLMs to enhance attention interpretability through natural language. “Attention Neural Baby Talk" [8917187] generates captions highlighting hazardous elements by aligning attention masks with descriptions, while DRAMA [malla2023drama] pairs videos with QA annotations explaining risk rationale. Large-scale VLMs like BLIP [li2022blip], Flamingo [alayrac2022flamingo], and LLaVA [10.5555/3666122.3667638] have enabled new approaches through vision-language pretraining and few-shot learning capabilities [xu2024drivegpt4interpretableendtoendautonomous, DriveVLM, zhou2024visionlanguagemodelsautonomous].
For driver attention specifically, LLada [zhou2025where] proposes joint modeling of attention maps and textual descriptions through their W3DA dataset ( 70k samples), predicting both where drivers look and why attention is allocated. GazeXplain [chen2024gazexplainlearningpredictnatural] generates natural language descriptions of gaze scanpaths for general visual attention.
2.3 Few-Shot Learning and Data-Efficient Adaptation
Few-shot learning addresses generalization from minimal examples. In semantic segmentation, PANet [wang2020panetfewshotimagesemantic] achieves 48.1% mIoU with 5 images per class, while few-shot object detection methods report 15-30% AP with 10-30 examples [kang2019fewshotobjectdetectionfeature, fan2020fewshotobjectdetectionattentionrpn, yan2024anomalysdfewshotmulticlassanomaly]. In generative modeling, DreamBooth [ruiz2023dreamboothfinetuningtexttoimage] personalizes Stable Diffusion using 3-5 images (CLIP-I 0.74), while UFC [kim2023universalfewshotlearningdense] achieves 87.3% accuracy with 30 examples versus 89.1% for fully-supervised baselines (10K+ examples).
Parameter-efficient fine-tuning enables adaptation with minimal updates. LoRA [hu2022lora] matches full fine-tuning performance while updating 0.01% of parameters, reducing GPT-3’s trainable parameters from 175B to 4.7M. For VLMs, LoRA adaptation maintains 95%+ performance with 10-20M trainable parameters [houlsby2019parameterefficienttransferlearningnlp]. Flamingo demonstrates in-context learning, improving VQAv2 from 49.2% (0-shot) to 63.1% (32-shot) [alayrac2022flamingo], though this approach shows high variance and struggles with structured outputs like spatial maps [dong2024surveyincontextlearning].
In autonomous driving, few-shot learning remains underexplored. Most attention systems train on hundreds of thousands of frames [palazzi2018predicting, xia2018predictingdriverattentioncritical, fang2019dada2000drivingaccidentpredicted].
Figure 2: Dataset curation pipeline. From BDD-A [10.1007/978-3-030-20873-8_42] videos to paired frames, GPT-4o captioning with fixed template, human verification, and final captions. Full details in Appendix.
3 Method
We present our few-samples dependent driver attention modeling framework that joints modeling of spatial driver attention prediction and structured natural language explanation. These two tasks depend on totally different kinds of spatial understanding, where captioning needs global semantic reasoning while gaze prediction requires localized spatial sensitivity. A shared cross-attention module tends to overfit one task and underfit another [Li_2024, fifty2021efficientlyidentifyingtaskgroupings]. To address this imbalance, we introduce a dual-pathway architecture (Figure 3) in which gaze prediction and caption generation are handled by separate modules while still leveraging shared visual features. A vision-language alignment mechanism further provides semantic supervision to spatial prediction, ensuring predicted attention regions correspond to meaningful visual content. This design enables effective joint learning despite data scarcity.
Figure 3: FSDAM architecture for joint gaze prediction and caption generation. A frozen vision-language backbone extracts spatial features $F_{t}$ and text embeddings $q_{\text{text}}$. The gaze pathway, left, upsamples $F_{t}$ to predict attention and gaze maps $\widehat{G}$. The caption pathway, right, uses cross-attention over $F_{t}$ to generate structured explanations.
3.1 Problem Formulation
Given a driving scene image $I\in\mathbb{R}^{H\times W\times 3}$, we jointly predict: (1) a spatial attention distribution $\widehat{G}\in\Delta^{S\times S}$ indicating where the driver looks, where $\Delta^{S\times S}$ denotes the probability simplex over an $S{\times}S$ grid with S=64S{=}64, and (2) a structured attention reasoning representation $\widehat{y}$ following the format:
$\widehat{y}=(C_{scene},C_{current},C_{next},C_{why})$ (1)
Here, $C_{{scene}}$ summarizes the global scene context, $C_{current}$ describes the driver’s present focus of attention, $C_{next}$ anticipates how the attention is likely to shift next (“will check the crosswalk”), and $C_{why}$ provides the causal rationale underlying this transition. Together, the four components form a structured attention reasoning representation that links spatial attention patterns with high-level semantic explanations, enabling the model to learn gaze behavior from single frames.
Our architecture builds upon LLaVA-Next-1.6 [liu2024llavanext], which comprises a CLIP-ViT-L/14 vision encoder [radford2021learning] extracting spatial features $F_{t}\in\mathbb{R}^{B\times 1024\times 24\times 24}$ and a Mistral-7B language model [jiang2023mistral7b] producing image-conditioned text embeddings $q_{\text{text}}\in\mathbb{R}^{B\times d_{\ell}}$ through multimodal fusion [alayrac2022flamingo, liu2024llavanext]. These frozen features feed into specialized pathways for gaze prediction and caption generation, with task-specific adaptation enabled by LoRA [hu2022lora]. We optimize 13.6M task-specific parameters (0.2% of the 7B backbone) while keeping pretrained components frozen to prevent overfitting.
3.2 Spatial Attention Prediction
The gaze pathway predicts where drivers look. A convolutional decoder upsamples $F_{t}$ from $24{\times}24$ to $64{\times}64$ resolution through successive convolution and bilinear upsampling layers [long2015fullyconvolutionalnetworkssemantic, ronneberger2015unetconvolutionalnetworksbiomedical], then applies spatial softmax to produce $\widehat{G}\in\Delta^{S\times S}$.
We supervise with forward KL divergence to encourage covering all human fixation regions:
$\mathcal{L}_{\text{KL}}=\text{KL}(G\|\widehat{G})=\sum_{i,j}G(i,j)\log\frac{G(i,j)}{\widehat{G}(i,j)}.$ (2)
This alone produces overly diffuse predictions under limited data. We augment with a blur-gap regularizer that applies Gaussian smoothing ($\sigma{=}1.0$) to $\widehat{G}$ to obtain $\widetilde{G}$, then penalizes cases where blurring doesn’t degrade prediction:
$\mathcal{L}_{\text{gaze}}=\mathcal{L}_{\text{KL}}+\lambda\cdot\max\big(0,\text{KL}(G\|\widetilde{G})-\mathcal{L}_{\text{KL}}+\epsilon\big),$ (3)
with $\lambda{=}0.3$ and $\epsilon{=}0.05$. This encourages confident spatial localization in low-data regimes [kümmerer2016deepgazeiireadingfixations, cornia2018predicting]. The gaze pathway receives gradients only from $\mathcal{L}_{\text{gaze}}$ during forward passes, with semantic supervision added through alignment (Section 3.4).
3.3 Attention-Grounded Caption Generation
Following prefix-tuning [li2021prefixtuning], we expand $q_{\text{text}}$ into $M$ queries via projection $W_{\text{cap}}$, then perform cross-attention [vaswani2017attention] over $F_{t}$:
$Q=W_{\text{cap}}(q_{\text{text}})\in\mathbb{R}^{B\times M\times d},\quad\text{CTX}=\text{CrossAttn}(Q,K,V).$ (4)
Mean pooling aggregates $M$ context vectors into $\bar{c}_{\text{cap}}$, which a Prefix Adapter $\Psi$ projects to visual prefix tokens $P=\Psi(\bar{c}_{\text{cap}})$ conditioning the decoder. We supervise with autoregressive cross-entropy:
$\mathcal{L}_{\text{cap}}=-\sum_{t=1}^{T}\log p_{\theta}(y_{t}\mid y_{<t},P,I),$ (5)
where $y$ contains scene description, current attention, anticipated shift, and causal reasoning. LoRA [hu2022lora] enables task-specific tuning while keeping most parameters frozen.
3.4 Vision-Language Alignment
To ensure predicted attention regions align with semantic content, we introduce a training-only alignment mechanism. A dedicated cross-attention block processes text queries $q_{\text{text}}$ to produce text-conditioned features $\bar{c}_{\text{gaze}}$. In parallel, we pool spatial features $F_{t}$ using predicted gaze $\widehat{G}_{\downarrow}$ (downsampled to match $F_{t}$ resolution) as soft weights:
$z_{\text{vis}}=\sum_{i,j}\widehat{G}_{\downarrow}(i,j)\,F_{t}[:,:,i,j].$ (6)
using the model’s own prediction maintains consistency between training and inference [chen2020simpleframeworkcontrastivelearning, grill2020bootstrap].
Both representations project into a shared 256-dimensional space via learned projections $P_{\text{vis}}$ and $P_{\text{txt}}$, producing $u^{\text{vis}}=P_{\text{vis}}(z_{\text{vis}})$ and $u^{\text{txt}}=P_{\text{txt}}(\bar{c}_{\text{gaze}})$. We align these with InfoNCE contrastive loss [oord2018representation]:
$\mathcal{L}_{\text{align}}=-\frac{1}{B}\sum_{i=1}^{B}\log\frac{\exp\big(\text{sim}(u_{i}^{\text{vis}},u_{i}^{\text{txt}})/\tau\big)}{\sum_{j=1}^{B}\exp\big(\text{sim}(u_{i}^{\text{vis}},u_{j}^{\text{txt}})/\tau\big)},$ (7)
where $\text{sim}(\cdot,\cdot)$ denotes cosine similarity and $\tau$ is a temperature parameter. Gradients pass through spatial pooling to the gaze decoder, pushing attention toward semantically relevant regions.
3.5 Training and Inference
We jointly optimize all three objectives through a weighted combination:
$\mathcal{L}=w_{g}\,\mathcal{L}_{\text{gaze}}+w_{c}\,\mathcal{L}_{\text{cap}}+w_{a}\,\mathcal{L}_{\text{align}},$ (8)
where $(w_{g},w_{c},w_{a})=(1.0,1.0,0.2)$ balance gaze prediction, caption generation, and vision-language alignment. The weights were chosen through a small-scale grid search on a validation split and were stable across experiments.
Training.
We train using AdamW optimizer with separate learning rates for LoRA and task modules. Training uses mixed precision, gradient accumulation for an effective batch of 16, and input size $336{\times}336$. Only LoRA and lightweight adapters are updated, totaling 13.6M trainable parameters ($\sim$0.2% of the backbone). Learning rate is adjusted via a ReduceLROnPlateau scheduler.
Inference.
During inference, only gaze and caption branches are active. The gaze decoder predicts $\widehat{G}$ from visual features $F_{t}$, while the caption pathway generates structured explanations through cross-attention and prefix adaptation. The alignment module is used only during training and adds no cost at inference.
4 Experimental Setup
Dataset Preparation.
We construct a gaze-language dataset from BDD-A dataset [10.1007/978-3-030-20873-8_42], which contains braking event videos selected from large-scale, crowd-sourced driving video data combined with compiled human gaze data from 6.5 seconds prior and 3.5 seconds after each braking event. To identify frames capturing meaningful attention transitions, we employ a KL-divergence-based selection algorithm [shlens2014noteskullbackleiblerdivergencelikelihood] that detects moments of maximum gaze distribution change between consecutive frames. Local peaks in the KL divergence curve mark anchor frames $t$ where attention shifts abruptly. For each anchor, we select a target frame $t+\Delta$ within $[3,18]$ frames that maximizes divergence from the anchor, capturing the strongest future attention transition. Following temporal sampling strategies from video action recognition [carreira2018quovadisactionrecognition], we filter clips shorter than 50 frames and retain the top - K = 2 anchor–target pairs per video for temporal diversity. Each resulting sample $(I_{t},I_{t+\Delta},G_{t},G_{t+\Delta})$ consists of two frames and their corresponding gaze maps. We generate structured captions via GPT-4o [openai2024gpt4technicalreport] with human verification (Figure 2), producing 90 training examples across eight diverse driving scenarios.
Implementation Details.
All experiments are implemented in PyTorch [paszke2019pytorchimperativestylehighperformance]. We build on the LLaVA-Next Mistral-7B backbone, which integrates a CLIP-based vision tower pretrained on large-scale image–text pairs. The vision–language backbone remains frozen during training, and only lightweight task-specific components (gaze decoder, prefix adapter, alignment head, and LoRA adapters) are updated. Training is performed on a single NVIDIA GH200 GPU with input resolution $336{\times}336$, batch size 4, and gradient accumulation over 4 steps (effective batch 16). We use the AdamW optimizer with learning rates of $1{\times}10^{-4}$ for LoRA and $2{\times}10^{-4}$ for task heads, a ReduceLROnPlateau scheduler (factor 0.5, patience 2), and mixed precision of bfloat16.
Evaluation Metrics
We evaluate our model using standard saliency and captioning metrics. For gaze prediction, we adopt CC, KL, SIM, AUC-J, AUC-B, and NSS following the official MIT Saliency Benchmark [bylinskii2017differentevaluationmetricstell]. For caption generation, we report BLEU, METEOR, ROUGE-L, CIDEr-R, and BERTScore using the COCO Caption Evaluation Toolkit [chen2015microsoftcococaptionsdata]. All metrics are computed using their standardized implementations to ensure fair comparison with prior work.
5 Results & Analysis
5.1 Baselines
We evaluate FSDAM on four driver attention datasets: BDD-A [10.1007/978-3-030-20873-8_42], DADA-2000 [fang2023dadadriverattentionprediction], DReyeVE [palazzi2018predicting], and W3D [zhou2025where], comparing against seven baselines: U²-Net [Qin_2020_PR], MINet [pang2020multiscaleinteractivenetworksalient], DBNet [9812524], DeepLabV3 [chen2017rethinkingatrousconvolutionsemantic], and LLada [zhou2025where]. These models represent current approaches in saliency and driver attention research. All baselines were trained on the complete BDD-A training set and tested on each dataset’s official test split. For LLada [zhou2025where], we report published results LLada was trained on the W3D dataset.
For caption prediction, we evaluate FSDAM on the W3D benchmark and compare it with models representing the current spectrum of driver attention approaches, from fully supervised to zero-shot and few-shot methods. Fully trained baselines including GazeXplain [chen2024gazexplainlearningpredictnatural] and LLada [zhou2025where] utilize the complete W3D [zhou2025where] training dataset (70k samples). Zero-shot and in-context models such as Qwen-VL [bai2025qwen25vltechnicalreport] and LLaVA [10.5555/3666122.3667638] are evaluated without task-specific fine-tuning, with Qwen-VL [bai2025qwen25vltechnicalreport] using in-context learning via fine-tuning prompts. Few-shot baselines combine gaze prediction models with language models—DeepGazeIIE [linardos2021deepgazeiiecalibratedprediction] + LLaVA [10.5555/3666122.3667638] and MLNet [8356626] + LLaVA [10.5555/3666122.3667638] employ a two-stage approach where gaze heatmaps are first generated then fed to the language model, both trained on the same 90 BDD-A samples as our method. FSDAM (Ours) uses an end-to-end architecture trained on 90 BDD-A samples and is evaluated on the W3D test set across three driving scenarios: normal driving, safety-critical situations, and traffic accidents, demonstrating the effectiveness of our unified gaze-language approach against both data-rich and data-efficient alternatives.
5.2 Quantitative Analysis
5.2.1 Spatial Attention Prediction
Table 1 reports quantitative results for driver attention prediction across four datasets. All reproduced models were trained on BDD-A for consistency. FSDAM provides a strong balance between sample efficiency and predictive accuracy. With only 90 training samples, it reaches performance that aligns with or exceeds models trained on much larger datasets, including W3D with $\sim$70K frames, DADA-2000 with 658K frames, and DR(eye)VE with  [^5] K frames.
On BDD-A[palazzi2018predicting], FSDAM obtains the lowest KL divergence at 1.13. This improves over LLada[zhou2025where] by 2.6% and over DBNet[9812524] by 13%. It matches LLada[zhou2025where] in CC at 0.60 and ranks second in NSS at 4.10, indicating better spatial attention prediction. Compared to the fully supervised MINet[pang2020multiscaleinteractivenetworksalient], FSDAM improves CC by 30% and reduces KL divergence by 89%, which highlights its effectiveness with limited data.
Our cross-dataset evaluation demonstrates the strong transferability of our approach. On DR(eye)VE[palazzi2018predicting], FSDAM reaches the lowest KL divergence at 0.77, improving over LLada[zhou2025where] by 26% and DBNet[9812524] by 57%. It also achieves the highest SIM at 0.54, a 20% gain over U²-Net[Qin_2020_PR]. This shows that vision language alignment helps capture domain stable attention cues that transfer to highway scenarios.
On DADA-2000[fang2023dadadriverattentionprediction], FSDAM maintains competitive performance while using far fewer samples, achieving the best KL divergence at 1.64 with a 10% gain over LLada[zhou2025where] and a 13% gain over DBNet[9812524]. On W3D, FSDAM matches DeepLabV3[chen2017rethinkingatrousconvolutionsemantic] in CC at 0.53 and beats LLada in KL divergence at 1.27 by 13%. It also achieves the second best SIM at 0.44 and NSS at 3.50.
These results show that extensive supervision is not required for strong driver attention modeling. FSDAM ranks first in KL divergence on three of the four datasets with an average 15% margin over the next best method. This establishes the proposed few shot vision language framework as a practical and scalable solution for new driving domains where large gaze annotation sets are not available.
Table 1: Quantitative comparison of driver attention prediction across four datasets. All reproduced models were trained on BDD-A dataset. Bold indicates best performance, underlined indicates second best. * denotes results reported from original papers.
Method BDD-A DADA-2000 DR(eye)VE W3D
[…]
U²-Net [Qin_2020_PR] 0.56 1.50 0.47 0.94 0.88 3.95 0.42 2.18 0.37 0.91 0.82 2.73 0.57 1.52 0.45 0.90 0.82 3.20 0.44 2.10 0.37 0.90 0.81 2.81
MINet [pang2020multiscaleinteractivenetworksalient] 0.46 10.3 0.16 0.89 0.82 3.67 0.32 10.45 0.14 0.82 0.73 2.02 0.44 8.87 0.36 0.84 0.80 2.65 0.35 10.51 0.14 0.83 0.77 2.54
DBNet [9812524] 0.57 1.30 0.41 0.95 0.91 4.41 0.41 1.89 0.29 0.93 0.85 3.08 0.48 1.79 0.29 0.91 0.85 3.71 0.47 1.77 0.33 0.93 0.86 3.50
DeepLabV3 [chen2017rethinkingatrousconvolutionsemantic] 0.47 9.62 0.21 0.97 0.75 2.56 0.42 10.26 0.19 0.96 0.73 2.23 0.67 8.78 0.47 0.91 0.84 2.74 0.53 9.70 0.32 0.93 0.81 2.35
LLada* [zhou2025where] 0.60 1.16 0.47 0.48 1.82 0.36 0.67 1.04 0.52 1.46 0.37 0.95 0.90 4.62
FSDAM (Ours) 0.60 1.13 0.43 0.96 0.91 4.10 0.44 1.64 0.33 0.92 0.86 2.95 0.61 0.77 0.54 0.91 0.84 4.04 0.53 1.27 0.44 0.91 0.87 3.50
5.2.2 Caption Prediction Analysis
Table 2 presents caption generation performance across three driving categories on the W3D dataset. Despite using only 90 training samples from BDD-A, our FSDAM achieves remarkable results that rival or exceed fully-trained baselines using the complete W3D dataset. In normal driving scenarios, FSDAM attains strong performance (BLEU: 0.42, METEOR: 0.37, CIDEr-R: 0.83) competitive with LLada (BLEU: 0.44, METEOR: 0.36, CIDEr-R: 0.96), which is trained on the full dataset. Notably, FSDAM demonstrates robust generalization to challenging scenarios, achieving competitive scores in safety-critical situations (BLEU: 0.35, METEOR: 0.33) and traffic accidents (BLEU: 0.33, METEOR: 0.35, CIDEr-R: 0.84). Compared to other few-shot baselines combining existing gaze models with LLaVA, FSDAM shows substantial improvements—achieving up to 3.2 times higher BLEU scores (0.42 vs. 0.13 for MLNet[pang2020multiscaleinteractivenetworksalient]+LLaVA in normal driving). Zero-shot approaches yield limited performance (BLEU << 0.21), underscoring the value of our few-shot learning framework, which effectively bridges the gap between zero-shot and full-data training regimes while preserving high performance efficiency.
Table 2: Comparison of captioning performance across driving scenarios on W3D dataset. Fully-trained baselines use the original W3D training data, Zero-shot and ICL models require no fine-tuning, and our FSDAM and few-shot baselines are trained on a 90-sample BDD-A subset. Higher is better.
Method Training Regime Normal Driving Safety-Critical Situation Traffic Accident
BLEU METEOR ROUGE CIDEr-R BLEU METEOR ROUGE CIDEr-R BLEU METEOR ROUGE CIDEr-R
Fully Trained on W3D
GazeXplain* [chen2024gazexplainlearningpredictnatural] Full-data (W3D) $\sim$70k samples] 0.31 0.30 0.22 0.42 0.19 0.29 0.37 0.55 0.17 0.20 0.44 0.66
LLada* [zhou2025where] Full-data (W3D) $\sim$70k samples] 0.44 0.36 0.58 0.96 0.44 0.38 0.59 1.23 0.38 0.32 0.52 1.00
Zero-shot and In-Context Models
Qwen-VL [bai2025qwen25vltechnicalreport] Zero-shot (no training) 0.10 0.19 0.28 0.34 0.19 0.21 0.29 0.13 0.08 0.21 0.29 0.12
LLaVA [10.5555/3666122.3667638] Zero-shot (no training) 0.12 0.14 0.23 0.35 0.13 0.19 0.11 0.10 0.17 0.26 0.19 0.13
Qwen-VL [bai2025qwen25vltechnicalreport] In-context learning (no fine-tuning) 0.13 0.18 0.22 0.36 0.21 0.17 0.30 0.23 0.12 0.24 0.33 0.15
Few-shot Learning (BDD-A 90 samples)
DeepGazeI[kümmerer2015deepgazeiboosting] + LLaVA Few-shot 0.12 0.23 0.28 0.14 0.13 0.22 0.30 0.18 0.15 0.21 0.31 0.17
DeepGazeIIE[linardos2021deepgazeiiecalibratedprediction] + LLaVA Few-shot 0.11 0.18 0.26 0.17 0.11 0.20 0.32 0.13 0.11 0.19 0.34 0.10
MLNet[8356626] + LLaVA Few-shot 0.13 0.19 0.27 0.31 0.26 0.20 0.32 0.12 0.13 0.18 0.33 0.28
FSDAM (Ours) Few-shot 0.42 0.37 0.48 0.83 0.35 0.33 0.46 0.47 0.33 0.35 0.34 0.84
5.3 Qualitative Results
Figure 4 presents qualitative comparisons against fully-supervised baselines U2-Net [Qin_2020_PR] and DeepLabV3 [chen2017rethinkingatrousconvolutionsemantic]. Row (a) demonstrates peripheral hazard detection where FSDAM produces a broader attention distribution extending toward the right-side pedestrian area, while baselines concentrate more tightly on the stop sign region. In the complex intersection scenario (row c) with multiple pedestrians and vehicles, FSDAM generates the most spatially extensive attention coverage across agents, while U2-Net produces more concentrated predictions and DeepLabV3 [chen2017rethinkingatrousconvolutionsemantic] exhibits multiple distinct hot spots, potentially missing critical areas. Row (e) illustrates left-turn navigation where FSDAM effectively captures distributed attention spanning both the lead vehicle and the turning trajectory, demonstrating the benefit of vision-language alignment in understanding maneuver semantics, while baselines show more localized attention patterns. Row (d) shows all methods successfully localizing attention to the lane-changing vehicle’s brake light area, with comparable spatial precision across approaches.
Row (b) presents a challenging scenario where ground truth exhibits distributed attention across both the lead vehicle and a right-side pedestrian. All methods concentrate primarily on the vehicle, missing the distributed pattern. Despite this shared challenge, FSDAM demonstrates strong performance across diverse driving scenarios, producing attention predictions with superior spatial coverage in multi-agent situations (rows a, c, e) and comparable precision in focused attention tasks (row d). These results validate that our vision-language alignment approach enables effective few-shot learning, achieving spatial and semantic coherence competitive with fully-supervised methods trained on substantially larger datasets
Figure 4: Qualitative comparison of driver attention prediction on BDD-A test scenes showing the input image and attention heatmaps from ground truth, FSDAM trained on 90 samples, U2-Net[Qin_2020_PR], and DeepLabV3[chen2017rethinkingatrousconvolutionsemantic].
Figure 5: Driving scene and model output showing the input frame, its ground truth gaze map, and the corresponding FSDAM prediction. The model produces attention heatmaps over key regions and generates structured attention reasoning.
5.4 Few-Shot Learning Analysis
Figure 1 presents the impact of varying support set sizes on gaze prediction performance across six metrics. FSDAM demonstrates efficient learning dynamics, with substantial improvements from 1-shot to 5-shot settings. Performance gains are most pronounced between 1-shot and 3-shot (e.g., NSS increases from 2.32 to 3.64, CC improves from 0.41 to 0.53), after which the improvements plateau. Notably, our 5-shot model achieves performance (CC: 0.58, SIM: 0.43, NSS: 4.09, KL: 1.17) that closely approaches the full-data baseline (CC: 0.60, SIM: 0.43, NSS: 4.10, KL: 1.13), recovering 96.7% of full-data CC performance and 99.8% of NSS performance with only five support samples. The KL divergence shows consistent reduction from 1.74 to 1.17 as support size increases, indicating improved distribution matching. This analysis validates FSDAM’s data efficiency: meaningful gaze predictions emerge from minimal supervision, with performance rapidly saturating after 3-5 examples, making our approach practical for real-world deployment where annotated driver attention data is scarce.
5.5 Ablations
We conduct systematic ablation experiments to assess each component’s contribution in FSDAM. For gaze prediction, we evaluate on the official BDD-A test set. For caption generation, we evaluate on 49 curated samples with structured explanations. We compare four variants: Gaze-Only trains only gaze prediction; Caption-Only trains only caption generation; Shared Cross-Attention uses a single cross-attention module for both tasks; and Full FSDAM uses task-specific cross-attention modules with vision–language alignment.
Figure 6 shows a consistent performance hierarchy across both tasks. The shared cross-attention variant provides marginal gains over single-task baselines but degrades gaze prediction (4% higher KL divergence), indicating negative transfer. In contrast, full FSDAM achieves substantial improvements: 31.9% KL reduction, 43.3% SIM improvement, and 66.7% NSS improvement for gaze prediction, alongside 32.4% ROUGE-L, 7.1% CIDEr-R, and 3.4% BERTScore improvements for caption generation, all relative to single-task baselines. This hierarchy (single-task << shared << dual-pathway) validates three design principles: joint training enables mutual benefits through shared visual representations, task-specific cross-attention modules prevent negative transfer, and dual pathways with vision–language alignment enable effective few-shot learning.
Figure 6: Ablation study comparing four architectural variants. The dual-pathway configuration in FSDAM achieves the largest percentage gains for both gaze prediction and caption generation.
6 Conclusions
We present FSDAM, a framework that achieves joint gaze prediction and natural language explanation from approximately 100 training examples. Through a dual-pathway architecture with vision-language alignment, we demonstrate competitive performance compared to fully-supervised methods, while introducing attention shift prediction. This paradigm also enables future implementation of a coupled sequential analysis framework, where forward predictions can be fed back into the system as guidance, functioning as policies or recommendations for performance adjustments. Our results validated this approach across multiple benchmarks (BDD-A, DR(eye)VE, DADA-2000, W3D), showing strong zero-shot generalization despite minimal supervision. The model generates contextually appropriate explanations across diverse driving scenarios, from normal conditions to safety-critical situations and accidents. This work demonstrated the feasibility of effective explainable attention modeling with manageable labeling. Our few-shot paradigm can facilitate practical deployment of visual attention prediction models in specialized domains where large-scale annotation may be infeasible. Future directions include temporal modeling through video inputs and explicit handling of distributed attention patterns to further improve anticipation accuracy in dynamic scenarios.